# Multi-perspective Feedback-attention Coupling Model for Continuous-time Dynamic Graphs

### v1

@@ Abstract
Recently, representation learning over graph networks has gained popularity, with various models showing promising results. Despite this, several challenges persist: 1) most methods are designed for static or discrete-time dynamic graphs; 2) existing continuous-time dynamic graph algorithms focus on a single evolving perspective; and 3) many continuous-time dynamic graph approaches necessitate numerous temporal neighbors to capture long-term dependencies. In response, this paper introduces the Multi-Perspective Feedback-Attention Coupling (MPFA) model. MPFA incorporates information from both evolving and raw perspectives, efficiently learning the interleaved dynamics of observed processes. The evolving perspective employs temporal self-attention to distinguish continuously evolving temporal neighbors for information aggregation. Through dynamic updates, this perspective can capture long-term dependencies using a small number of temporal neighbors. Meanwhile, the raw perspective utilizes a feedback attention module with growth characteristic coefficients to aggregate raw neighborhood information. Experimental results on a self-organizing dataset and seven public datasets validate the efficacy and competitiveness of our proposed model.
@@
Continuous-time dynamic graphs, Multi-perspective, Feedback attention, Evolving and raw perspectives, Representation learning
publicationid: pubid: $0$–$00$ 2021 IEEE
Figure 1: A toy example illustrates a continuous-time dynamic graph with multiple complex dynamics, where $t_{0}$-$t_{3}$ represent timestamps. The evolution of the graph from $t_{0}$ to $t_{3}$ captures the entire process, revealing that new interactions between nodes are established at each timestamp. At $t_{1}$, a new node joins the network and interacts with another node (red edge). At $t_{2}$, co-occurrence events occur (green edges), and at the final timestamp, the connection between nodes 1 and 6 is re-established (blue edge). Conversely, a discrete-time dynamic graph can be formed by extracting snapshots (e.g., snapshots $t_{0}$ and $t_{2}$) from the temporal network. If only the final state of the graph evolution is considered, a static graph can be obtained.
@@ I Introduction
Representation learning has illuminated the analysis of graph-structured data, leveraging its powerful capability to encode structures and properties of graph networks. Typically, it acquires node embedding representations by transforming high-dimensional and non-Euclidean graphs into lower-dimensional vector embeddings [1]. These node embeddings are then employed in various graph analysis tasks, including graph classification, edge prediction, or node classification [2, 3]. While existing representation learning methods have demonstrated promising performance across numerous domains [4, 5], several challenges remian. The first issue with the majority of graph representation learning methods is that they primarily focus on static graph networks or discrete-time dynamic graphs[6, 7, 8]. In reality, graph data in most domains (e.g., natural language processing, social networks, recommendation systems) exhibit complex dynamics, meaning that the network structures and properties constantly change over time (In Figure 1, we illustrate a straightforward process depicting changes in a dynamic network). The second problem is that existing continuous-time dynamic graph algorithms only consider the evolving perspective, ignoring other aspects of the graph. Finally, most continuous-time dynamic graph approaches require a large number of temporal neighbors to capture long-term dependencies, which can be computationally expensive. Models for continuous-time dynamic graph representation learning need to support the addition of nodes and edges, as well as feature changes, and generate node embeddings at any time. While the usage of static graph methods (e.g., GAT [8] and GraphSAGE [9]) on temporal networks is possible in some cases, taking no account of the temporal information can severely weaken the ability to capture the crucial insights of networks. Therefore, there is a need for the development of more and more effective models for these dynamically changing data.
@@
Existing techniques for learning representations on dynamic graphs can be divided into two folds: discrete-time dynamic graphs and continuous-time dynamic graphs. Discrete-time dynamic graphs are represented as a sequence of network snapshots at specific time intervals [10, 11, 12]. Since the network is represented in snapshots, the static graph analysis method can be used for each snapshot with a fixed topology. The repeated static method and the time information learning between the snapshot embeddings can collectively give insight into the graph’s dynamics. The existing methods of discrete dynamic graphs typically include stacked structures, integrated structures, graph auto-encoders, and generative models [13, 14, 15, 16, 17, 18]. However, such dynamic graph representation makes a model sensitive to the snapshot sampling interval. When the sampling interval is too large, much information in the network will be lost, leading to the model only capturing limited evolutionary information. Recently, several methods have been proposed for supporting the continuous-time settings[19, 20, 21]. These methods view dynamic graphs as a stream of events occurring in chronological order, so their typical processing methods follow an event-level paradigm. Examples include temporal random walk-based methods[22, 23], RNN-based methods[24, 25], attention-based methods[26], and transformer-based methods[27, 28]. We find that these methods need to look up a large number of historical interactions in order to obtain long-term dependencies[29, 30, 28]. This leads to increased computational and time costs. When the acquired historical interactions are short, the performance of the model degrades. In addition, these methods focus only on a single evolving perspective. Although memory-based methods can use shorter historical interactions to acquire long-term dependencies, they are also limited by the singularity of the perspective, which prevents the model from learning the nature of the raw interactions and the growth characteristics from the raw interactions to the current state.
@@
In this paper, we propose corresponding solutions for each of the above challenges: For the first challenge, our proposed MPFA can effectively complement the continuous-time dynamic graph model. For the second challenge, due to the time-varying nature of continuous-time dynamic graphs, when a model aggregates the temporal neighbor information of a node, these neighbors may no longer be in the interaction initial state (they may have interacted with other nodes multiple times). In many application scenarios, only considering evolutionary interactions is often insufficient and may lead to information redundancy or inference errors. (For example, in the shopping system, after a person has purchased a mobile phone, his purchase of a mobile phone protective case is generally related to the purchase of a mobile phone and should not be affected too much by other events). Therefore, we propose to preserve the raw state to improve the performance and generalization of a model. First, we design an evolving perspective that can capture the real-time dynamics of the nodes at any time. All participating nodes in this perspective contain the latest evolutionary content of the network. To distinguish the importance of different neighbors, we use temporal self-attention to discriminatively aggregate the neighbor information. Second, we design a raw perspective. The raw interaction represents the past facts and reveals the essence of events. When the network returns to observe these events after learning, it will have deeper insights. We achieve the information fusion of the raw interactions by using the feedback attention coefficients with model growth characteristics. Finally, the two perspectives learn interactively, instructing each other to obtain the final target node embedding. For the third challenge, to achieve the property of capturing long-term dependencies using a small number of temporal neighbors, we perform a dynamic update of the evolving information in the evolving perspective immediately after each event. This allows a form of compressed coding to record the historical events for each node. Overall, the main contributions of this paper are as follows:
@@
  • We propose a novel inductive model called MPFA, working on continuous-time dynamic graphs described as an ordered sequence of timepoint events.
@@
  • We make the first attempt to explore continuous-time dynamic graphs from raw and evolving perspectives and design different attentions for them. The two views enable our model perform better in inductive prediction tasks as well as in dynamic classification tasks.
@@
  • We design a feedback attention coefficient with model growth characteristics to learn the neighborhood information closely related to the target node with deeper insights. In addition, dynamically updating evolving information can be used to capture long-term dependencies using a small number of temporal neighbors.
@@ II Related work
In this section, we briefly introduce the related work on static and discrete-time dynamic graphs and then elaborate on the existing work on continuous-time dynamic graphs.
@@
Figure 2: The Interaction between raw and evolving information (i.e., transmission of information on a timeline). The model learns node embedding based on the two informations and predicts whether event (1, 6) occurs at time $t$. For node 1 (similar to node 6), the dotted arrow represents the influence of evolving information on the raw information, and vice versa. The blue and red arrows mean the effect of the raw and evolving neighborhood on the current event; The effect of the node’s growth information on the raw ation is indicated by the orange dotted arrows. The purple dotted arrows represent that the evolving neighborhood has an impact on its own raw information.
@@
Continuous-time dynamic graphs: Processing continuous-time dynamic graphs is a challenging task, and only recently, some approaches have been proposed [38]. CTDNE[22] modifies the path sampling method of DeepWalk to make the path meet the actual occurrence sequence according to time, thereby making the sampling path more reasonable. Based on LSTM, the architecture of DyGNN[19] includes two components: An update module updating the embedding of the nodes involved in an event; A propagation module responsible for propagating the updated state to the node neighbors. JODIE [24] employs two RNNs to design a coupled recurrent neural network for node embedding, containing three modules: update, projection, and future interaction prediction. Some methods are based on variants of random walks[20, 23], which integrate continuous-time information by constraining on transition probability matrix. TGAT[26] takes the self-attention mechanism as a building block and develops a novel functional temporal encoding technique according to the classical Bochner theorem of harmonic analysis. DYREP[21] considers two dynamic processes on the temporal network: the Association Process and the Communication Process. The former represents changes in network topology, and the latter denotes changes in dynamics on the network. TGN [25] designed a node memory module to store the long-term dependencies of nodes. CAWN [29] initially generates a multitude of causal anonymous walks for each node. Subsequently, it encodes each walk using a recurrent neural network and combines these walks to attain the ultimate node representation. EdgeBank [30] is a purely memory-based method with no trainable parameters for transductive dynamic prediction tasks. TCL [27] uses a breadth-first search algorithm to generate interaction sequences of nodes and a graph transformer to obtain node embedding representations. DyGFormer[28] uses Transformer as the encoder. Neighbor co-occurrence coding as well as patching techniques are used in the sequence processing scheme. GraphMixer [39] uses a time encoding that requires no training and uses MLP-Mixer as the encoder to obtain node embedding representations.
@@
Figure 3: The overall architecture of MPFA. $t_{p_{1}}-t_{p_{6}}$ denote the interaction time before time $t$. The evolving perspective (bottom) aggregates information from constantly evolving neighborhood by using a temporal self-attention module; For the raw perspective (top), the feedback attentions with growth characteristics calculated by the Feedback Coefficient module are utilized to aggregate the raw neighborhood information; In the above two processes, the Raw and Evolving Information Preserved modules play a critical role. Then the acquired evolving and raw features are coupled to teach each other through the Attention Coupling module for final temporal graph embedding. After completing the learning of the two perspectives (gray arrow process), MPFA performs the update of EIP-RIP (red arrow process).
III Definations
@@ III-A Continuous-time dynamic graphs
A continuous-time dynamic graph is described as a sequence of time point events, where each event involves two nodes at a certain time. Formally, it can be represented as $G(T)=(V(T),E(T))=[x({t_{0}}),\;x({t_{1}}),\;x({t_{2}}),\>...]$. Given any time $t$, it can be denoted as $G(t)=(V(t),E(t))=[x({t_{0}}),x({t_{1}}),...,\>x({t})]$, where $V(t)$ and $E(t)$ are sets of nodes and edges, respectively. $x(t)=(v_{i},v_{j},e_{i,j}(t))$ is an event happened at time $t$, where $v_{i}$ and $v_{j}$ mean the nodes, $i$ and $j$ are the node numbers. $e_{i,j}(t)$ denotes the edge feature. It should be noted that a pair of nodes $(v_{i},v_{j})$ may build multiple interactions at different time points.
III-B Continuous-time dynamic graph representation learning
@@
The goal of Continuous-time dynamic graph representation learning is to learn node representations and use them for downstream tasks (e.g., dynamic node classification, and future link prediction). Given a temporal graph $G$, temporal graph embedding aims to seek a mapping function $f:V\longrightarrow\mathbb{R}^{d}$, where $d$ is the dimension of node representation. At the time $t$, the temporal graph embedding can be denoted by $Z(t)=({z_{1}}(t),...,{z_{n}}(t))$.
III-C Raw and evolving information preserved
According to the evolution direction of the network in time, we can divide the temporal events of a node into two states, the raw and evolving state. As shown in Figure 2, the raw state maintains the essence of an event at that time, and the evolving state is a description of a node’s ongoing change.
Evolving information preserved: Different from conventional approaches, we design a non-parameterized vector for each node to preserve the evolving information instead of directly parameterizing the node embedding. The evolving information of node $v_{i}$ is denoted as ${h_{i}}(t)$ at time $t$, representing all events of the node that occurred before time t in a compressed encoding form. This encoding captures long-term dependencies while utilizing only a small number of temporal neighbors. During the network evolving, it is updated after each event. When an event $x(t)$ happens between node $v_{i}$ and $v_{j}$, we first obatin the current event information ${X_{i,j}(t)}=W_{evol}[{z_{i}}({t^{-}})||{e_{i,j}}(t^{-})||{z_{j}}({t^{-}})||\Phi(\Delta t)]$, and then update the evolving information of $v_{i}$ as follows:
${h_{i}}(t)=update({X_{i,j}(t)},\;{h_{i}}({t^{-}})),$ (1)
where $\Phi(\cdot)$ denotes a generic time embedding [40, 26], and ||||| | represents a connection symbol. ${t^{-}}$ is the last interaction time of node $v_{i}$. $W_{evol}$ is the parameter used to scale the dimension. $z_{i}(t^{-})$ and $z_{j}(t^{-})$ denote the last embedding representation of nodes $v_{i}$ and $v_{j}$, respectively. Update(.) means a learnable function, which can be GRU or RNN. When a new node appears in the network, its evolving information will be initialized to a vector with zero values.
Raw information preserved: Due to the time-varying nature of continuous-time dynamic networks, each node will change (e.g., interacting with other nodes) over time. Thus, the nodes would no longer maintain the original interactive state. It is essential to keep the initial event state between node $v_{i}$ and $v_{j}$ unchanged so that the model can better understand the nature of historical events. It can be considered a growth process from the raw state to the evolving state. Given the raw events $[(v_{i},v_{1},e_{i,1}(t_{1})),...,(v_{i},v_{j},e_{i,j}(t_{j}))]$ of $v_{i}$, their raw event information is represented as $[{r_{1}}({t_{1}}),...,{r_{j}}({t_{j}})]$. How to design the raw information of a node is worth exploring, which will also be one of the directions of our future work. Here, for convenience, We denote the raw information as ${r_{j}(t_{j})}=W_{raw}[{z_{j}}({t^{-}_{j}})||{e_{i,j}}(t_{j}^{-})||{z_{i}}({t^{-}_{j}})||\Phi(\Delta t_{j})]$.
IV Representation learning on raw and evolving perspectives
Given the problem definitions above, we describe our proposed model in this section. We first illustrate the general framework of the MPFA in Figure 3 and then elaborate on its composition. MPFA, a multi-perspective feedback-attention coupling model, can project the temporal graph $G(t)$ at time t into low-dimensional space to generate node embeddings. The evolving perspective of MPFA consists of a temporal self-attention module aggregating constantly evolving temporal neighborhood information. In this evolving perspective, dynamically updated evolutionary information can capture long-term dependencies using fewer temporal neighbors. The raw perspective is an feedback attention for generating the aggregation of the raw neighborhood. To implement the feedback attention, the feedback attention coefficients with model growth characteristics are computed by a feedback module. Then, the two perspectives learn collaboratively to enhance the generalization and prediction capabilities of the model. More specific details of each module are discussed in the following subsections.

### v2

@@ Abstract
Representation learning over graph networks has recently gained popularity, with many models showing promising results. However, several challenges remain: 1) most methods are designed for static or discrete-time dynamic graphs; 2) existing continuous-time dynamic graph algorithms focus on a single evolving perspective; and 3) many continuous-time dynamic graph approaches necessitate numerous temporal neighbors to capture long-term dependencies. In response, this paper introduces a Multi-Perspective Feedback-Attention Coupling (MPFA) model. MPFA incorporates information from both evolving and original perspectives to effectively learn the complex dynamics of dynamic graph evolution processes. The evolving perspective considers the current state of historical interaction events of nodes and uses a temporal attention module to aggregate current state information. This perspective also makes it possible to capture long-term dependencies of nodes using a small number of temporal neighbors. Meanwhile, the original perspective utilizes a feedback attention module with growth characteristic coefficients to aggregate the original state information of node interactions. Experimental results on one dataset organized by ourselves and seven public datasets validate the effectiveness and competitiveness of our proposed model.
@@
Continuous-time dynamic graphs, Multi-perspective, Feedback-attention, Evolving and original perspectives, Representation learning
Figure 1: A simple example illustrates a continuous-time dynamic graph with multiple complex dynamics, where $t_{0}$-$t_{3}$ represent timestamps. The evolution of the graph from $t_{0}$ to $t_{3}$ captures the entire process, revealing that new interactions between nodes are established at each timestamp. At $t_{1}$, a new node joins the network and interacts with another node (red edge). At $t_{2}$, co-occurrence events occur (green edges), and at the final timestamp, the connection between nodes 1 and 6 is re-established (blue edge). Conversely, a discrete-time dynamic graph can be formed by extracting snapshots (e.g., snapshots $t_{0}$ and $t_{2}$) from the temporal network. If only the final state of the graph evolution is considered, a static graph can be obtained.
@@ I Introduction
Representation learning has illuminated the analysis of graph-structured data, leveraging its powerful capability to encode structures and properties of graph networks. Typically, it acquires node embedding representations by transforming high-dimensional and non-Euclidean graphs into lower-dimensional vector embeddings [1]. These node embeddings are then employed in various graph analysis tasks, including graph classification, edge prediction, or node classification [2, 3]. While existing representation learning methods have demonstrated promising performance across numerous domains [4, 5], several challenges remain. The first problem with most graph representation learning methods is that they focus primarily on static graph networks or discrete-time dynamic graphs[6, 7, 8]. In reality, graph data in most domains (e.g., natural language processing, social networks, recommendation systems) exhibit complex dynamics, i.e., the network structures and properties are constantly changing over time (in Figure 1, we illustrate a simple process showing changes in a dynamic network). The second problem is that existing continuous-time dynamic graph algorithms consider only the evolving perspective, ignoring other aspects of the graph. Finally, most continuous-time dynamic graph approaches require a large number of temporal neighbors to capture long-term dependencies, which can be computationally expensive. Models for continuous-time dynamic graph representation learning need to support the addition of nodes and edges, as well as feature changes, and generate node embeddings at any time. While the usage of static graph methods (e.g., GAT [8] and GraphSAGE [9]) on temporal networks is possible in some cases, taking no account of the temporal information can severely weaken the ability to capture the crucial insights of networks. Therefore, there is a need for the development of more and more effective models for these dynamically changing data.
@@
Existing techniques for learning representations on dynamic graphs can be divided into two types: discrete-time dynamic graphs and continuous-time dynamic graphs. Discrete-time dynamic graphs are represented as a sequence of network snapshots at certain time intervals [10, 11, 12]. Since the network is represented in snapshots, the static graph analysis method can be used for each snapshot with a fixed topology. The iterative static method and the temporal learning between the snapshot embeddings can together provide insight into the dynamics of the graph. The existing methods of discrete dynamic graphs typically include stacked structures, integrated structures, graph auto-encoders, and generative models [13, 14, 15, 16, 17, 18]. However, such a dynamic graph representation makes a model sensitive to the snapshot sampling interval. If the sampling interval is too large, much information in the network will be lost, resulting in the model capturing only limited evolutionary information. Recently, several methods have been proposed to support continuous-time settings[19, 20, 21]. These methods view dynamic graphs as a stream of events occurring in time order, so their typical processing methods follow an event-level paradigm. Examples include temporal random walk-based methods[22, 23], RNN-based methods[24, 25], attention-based methods[26], and transformer-based methods[27, 28]. We find that these methods need to look up a large number of historical interactions to obtain long-term dependencies[29, 30, 28]. This leads to increased computational and time costs. If the acquired historical interactions are short, the performance of the model degrades. In addition, these methods only focus on a single evolving perspective. Although memory-based methods can use shorter historical interactions to acquire long-term dependencies, they are also limited by the singularity of the perspective, which prevents the model from learning the nature of the original interactions and the growth characteristics from the original interactions to the current state.
@@
In this paper, we propose appropriate solutions for each of the above challenges. For the first challenge, our proposed MPFA can effectively complement the continuous-time dynamic graph model. For the second challenge, due to the time-varying nature of continuous-time dynamic graphs, when a model aggregates the temporal neighbor information of a node, these neighbors may no longer be in the initial interaction state (they may have interacted with other nodes multiple times). In many application scenarios, considering only evolutionary interactions is often insufficient and may lead to information redundancy or inference errors. (For example, in a shopping system, after a person buys a cell phone, his or her purchase of a cell phone case is generally related to the purchase of a cell phone and should not be influenced too much by other events). Therefore, we model dynamic graphs from both an evolving and an original perspective. The evolving perspective captures the current state information of the network, which can be used to describe the real-time evolutionary characteristics of the network. The original perspective represents past facts, revealing the essence of events. The two perspectives interact through mutual guidance, allowing the acquisition of node embeddings with more general properties. For the third challenge, to achieve the property of capturing long-term dependencies using a small number of temporal neighbors, we perform a dynamic update of the current state information in the evolving perspective immediately after each event. This allows a form of compressed coding to record the historical events for each node. Overall, the main contributions of this paper are as follows:
@@
  • We propose a novel inductive model called MPFA, which works on continuous-time dynamic graphs described as an ordered sequence of time-point events.
@@
  • We make the first attempt to explore continuous-time dynamic graphs from the original and evolving perspectives, and to design different attentions for them. The two views allow our model to perform better in both inductive link prediction and dynamic classification tasks.
@@
  • We design a feedback attention coefficient with model growth characteristics to learn the original state information closely related to the target node with deeper insights. In addition, dynamically updating evolving state information can be used to capture long-term dependencies using a small number of temporal neighbors.
@@ II Related work
In this section, we briefly introduce the related work on static and discrete-time dynamic graphs and then elaborate on continuous-time dynamic graphs.
@@
Figure 2: The Original and current network states from two perspectives. The model learns embedding representations for nodes $v_{i}$ and $v_{6}$ at time $t$ and predicts the occurrence of the event $x(t)=(v_{i},v_{6},e_{1,6}(t))$. For node $v_{i}$ (and similarly for node $v_{6}$), the orange dashed arrow represents feedback from the current state of node $v_{i}$ to the original state of the interacting node. The purple dashed arrow represents feedback from the current states of node $v_{i}$’s neighbors to their original states. The blue and red solid arrows show the influence of original and current state information on the embedding representation of node $v_{i}$, respectively.
@@
Continuous-time dynamic graphs: Processing continuous-time dynamic graphs is a challenging task, and only recently have some approaches been proposed [38]. CTDNE[22] modifies the path sampling method of DeepWalk to make the path match the actual sequence of occurrences according to time, thus making the sampling path more reasonable. Based on LSTM, the architecture of DyGNN[19] includes two components: An update module, which updates the embedding of the nodes involved in an event; A propagation module, which is responsible for propagating the updated state to the nodes’ neighbors. JODIE [24] uses two RNNs to design a coupled recurrent neural network for node embedding that includes three modules: update, projection, and future interaction prediction. Some methods are based on variants of random walks[20, 23] that integrate continuous-time information by constraining the transition probability matrix. TGAT[26] takes the self-attention mechanism as a building block and develops a novel functional temporal coding technique according to the classical Bochner theorem of harmonic analysis. DYREP[21] considers two dynamic processes in the temporal network: the association process and the communication process. The former represents changes in the network topology, and the latter represents changes in the dynamics of the network. TGN [25] designed a node memory module to store the long-term dependencies of nodes. CAWN [29] first generates a set of causally anonymous walks for each node. It then encodes each walk using a recurrent neural network and combines these walks to obtain the final node representation. EdgeBank [30] is a purely memory-based method with no trainable parameters for transductive dynamic prediction tasks. TCL [27] uses a breadth-first search algorithm to generate interaction sequences of nodes and a graph transformer to obtain node embedding representations. DyGFormer[28] uses the transformer as encoder. Neighbor co-occurrence coding and patching techniques are used in the sequence processing scheme. GraphMixer [39] uses a time encoding that requires no training and uses MLP-Mixer as the encoder to obtain node embedding representations.
@@
Figure 3: The overall architecture of MPFA. $t_{p_{1}}-t_{p_{6}}$ denote the interaction time before time $t$. The evolving perspective (bottom) aggregates the current state information of historical events by using a temporal self-attention module; For the original perspective (top), the feedback attentions with growth characteristics calculated by the feedback coefficient module are utilized to aggregate the original state information; In the above two processes, the original and evolving information preserved modules play a critical role. Then, the acquired evolving and original features are coupled to teach each other by the attention coupling module for the final temporal graph embedding. After completing the learning of the two perspectives (gray arrow process), MPFA performs the updating of the ESP-OSP (red arrow process).
III Preliminaries
@@ III-A Continuous-time dynamic graphs
We describe continuous-time dynamic graphs as a sequence of interaction events occurring in chronological order. Formally, it can be represented as $G\left(T\right)=\left(V\left(T\right),E\left(T\right)\right)=\left[x\left(t_{0}\right),\;x\left(t_{1}\right),\;x\left(t_{2}\right),\>\ldots\right]$. Given time $t$, it can be expressed as $\ G\left(t\right)=\left(V\left(t\right),E\left(t\right)\right)=\left[x\left(t_{0}\right),x\left(t_{1}\right),\ldots,\>x\left(t\right)\right]$, where $V\left(t\right)$ and $E\left(t\right)$ are the sets of nodes and edges at time $t$ respectively. $x\left(t\right)=\left(v_{i},v_{j},e_{i,j}\left(t\right)\right)$ denotes an event occurring at time $t$, where $v_{i}$ and $v_{j}$ represent nodes, and $i$ and $j$ are node indices. $e_{i,j}(t)$ represents the temporal edge feature.
III-B Continuous-time dynamic graph embedding learning
@@
The goal of continuous-time dynamic graph embedding learning is to acquire node feature representations and utilize them for downstream tasks such as dynamic node classification or future link prediction. Given a dynamic graph $G(t)$ at time $t$, its embedding learning aims to seek a mapping function $f:V\longrightarrow R^{d}$, where $d$ is the dimension of node representations. After the mapping function $f$, $G(t)$ can be represented as $\ Z\left(t\right)=\left(z_{0}\left(t\right),\ z_{1}\left(t\right),...\right)$.
III-C Original state and current state
We analyze the evolution of dynamic graphs from two perspectives: original and evolving. The evolving perspective is employed to capture the current state of nodes, while the original perspective is utilized to capture the original state of nodes (these two states are shown in Figure 2). Specifically, given a node $v_{i}$ at time $t$, its historical interaction at time $t_{1}$($t_{1}<t$) is denoted as $(v_{i},v_{1},e_{i,1}\left(t_{1}\right))$, where the states of nodes $v_{i}$ and $v_{1}$ at $t_{1}$ are are represented as $r_{i}\left(t_{1}\right)$ and $r_{1}\left(t_{1}\right)$, respectively. During the interval from $t_{1}$ to $t$, due to potential changes in nodes $v_{i}$ or $v_{1}$ (e.g., interactions with other nodes), the states of nodes $v_{i}$ and $v_{1}$ evolve into ${\ h}_{i}\left(t\right)$ and ${\ h}_{1}\left(t\right)$, respectively. We refer to $r_{i}\left(t_{1}\right)$ as the original state of node $v_{i}$ and $h_{i}\left(t\right)$ as the current state of node $v_{i}$. The original state retains the essence of the interaction events, while the current state records the real-time changes of the nodes.
III-D Current state preserved
In this section, we design a vector with an initial value of zero for each node, aiming to preserve the current state information of the nodes. The current state of node $v_{i}$ at time $t$ is represented as $h_{i}\left(t\right)$, which compresses and encodes all the events that occurred before time $t$ for that node. This encoding method requires only a small number of temporal neighbors to capture the long-term dependencies of the node. During the evolution of the network, whenever an event occurs for node $v_{i}$, its current state is updated. This process involves incorporating the event $x\left(t^{-}\right)$, occurring at node $v_{i}$ at time $t^{-}$, into $h_{i}\left(t^{-}\right)$ to derive $h_{i}\left(t\right)$,
$X_{i,j}(t^{-})=W_{evol}[z_{i}(t^{-})||e_{i,j}(t^{-})||z_{j}(t^{-})||\Phi\left(t-t^{-}\right)],$ (1)
$h_{i}\left(t\right)=update\left(X_{i,j}\left(t^{-}\right),\;h_{i}\left(t^{-}\right)\right),$ (2)
where $W_{evol}$ is a trainable parameter, $\Phi\left(\cdot\right)$ represents a generic time encoding, and ||||| | denotes the concatenation operator. $t^{-}$ is the most recent interaction time of node $v_{i}$. $z_{i}\left(t^{-}\right)$ and $z_{j}\left(t^{-}\right)$ represent the embedding representations of nodes $v_{i}$ and $v_{j}$ at time $t^{-}$, respectively. $update()$ denotes a learnable update function, which can be GRU or RNN.
III-E Original state preserved
Due to the time-varying nature of continuous-time dynamic networks, each node undergoes changes over time (such as interactions with other nodes), and thus no longer maintains its initial interaction state. However, the original state of a node represents the essence of interaction events. Therefore, we propose to preserve the original state of node $v_{i}$ at the time of interaction with $v_{j}$. Specifically, given an interaction $\left(v_{i},v_{j},e_{i,j}\left(t_{j}\right)\right)$ of node $v_{i}$ before time $t$, the calculation of the node $v_{i}$’s original state information $r_{i}\left(t_{j}\right)$ is as follows,
$r_{i}(t_{j})=W_{raw}[z_{i}(t_{j}^{-})||e_{i,j}(t_{j}^{-})||z_{j}(t_{j}^{-})||\Phi(t_{j}-t_{j}^{-})],$ (3)
where $W_{raw}$ represents the trainable parameters. For the original state, we allocate a temporary storage unit, which is discarded whenever the original state of a node’s neighbor is no longer in use.
IV Multi-perspective feedback-attention coupling learning
Given the above symbolic representations and definitions, this section introduces the proposed model. The general framework of MPFA is shown in Figure 3. MPFA is a multi-perspective feedback-attention coupling model that learns the evolution process of dynamic graphs from two perspectives. The evolving perspective considers the current state of historical interaction events of nodes and uses a temporal attention module to aggregate the current state information of these events. In this perspective, dynamically updating the current state information allows to capture long-term dependencies of nodes with fewer historical events. The original perspective uses feedback attention with a growth property to aggregate the original state information of node interactions. To implement feedback attention, the feedback module in the model computes feedback attention coefficients with growth properties. Finally, the two perspectives learn cooperatively to improve the generalization and prediction capabilities of the model.